Use the normal distribution of fish lengths for which the mean is 10 inches and the standard deviation is 2 inches. Assume the variable x is normally distributed.

a)What percent of the fish are longer than 14 inches?

b)If 400 fish are randomly selected about how many would you expect to be shorter than 9 inches?
Answer #1

a)

for normal distribution z score =(X-μ)/σ
here mean=       μ= 10
std deviation   =σ= 2.0000
probability = P(X>14) = P(Z>2)= 1-P(Z<2)= 1-0.9772= 0.0228~2.28%

b)

probability = P(X<9) = P(Z<-0.5)= 0.3085

hence expected number of fish which are shorter than 9 inches =np=400*0.3085=123.4